# Accelerating Diffusion Transformer via Error-Optimized Cache

### v2

@@ Abstract.
Diffusion Transformer (DiT) is a crucial method for content generation. However, it needs a lot of time to sample. Many studies have attempted to use caching to reduce the time consumption of sampling. Existing caching methods accelerate generation by reusing DiT features from the previous time step and skipping calculations in the next, but they tend to locate and cache low-error modules without focusing on reducing caching-induced errors, resulting in a sharp decline in generated content quality when increasing caching intensity. To solve this problem, we propose the Error-Optimized Cache (EOC). This method introduces three key improvements: (1) Prior knowledge extraction: Extract and process the caching differences; (2) A judgment method for cache optimization: Determine whether certain caching steps need to be optimized; (3) Cache optimization: reduce caching errors. Experiments show that this algorithm significantly reduces the error accumulation caused by caching, especially excessive caching. On the ImageNet dataset, without substantially increasing the computational load, this method improves the FID↓ of the generated images when the rule-based model FORA has a caching level of 75%, 50%, and 25%, and the training-based model Learning-to-cache has a caching level of 22%. Specifically, the FID↓ values change from 30.454 to 21.690 (28.8%), from 6.857 to 5.821 (15.1%), from 3.870 to 3.692 (4.6%), and from 3.539 to 3.451 (2.5%) respectively.
@@
ccs: Computing methodologies Computer vision
@@ 1. Introduction
Content generation aims to automatically produce text (Achiam et al., 2023), images (Xu et al., 2024), videos (Wang et al., 2024), and other forms of content based on existing data or context. It is widely used in intelligent customer service dialogue generation, news summarization, and advertising copy creation. However, current generation methods often rely on complex models, resulting in cumbersome and time-consuming computational processes that hinder rapid deployment and flexible application in real-world business contexts.
@@ 2.1. Traditional Generation Acceleration
Model Pruning And Quantization. Pruning reduces model complexity and computational cost by removing less critical weights, neurons, or layers while maintaining performance. It involves unstructured pruning (Dong et al., 2017; Lee et al., 2019), which masks individual parameters, and structural pruning (Liu et al., 2021), which removes larger components like layers or filters. In diffusion models, techniques like LD-Pruner (Castells et al., 2024) and DiffPruning (Guo et al., 2020) enhance efficiency by applying structured pruning and using informative gradients to reduce unnecessary steps. Quantization lowers the precision of weights and activations to smaller bit formats, reducing model size and speeding up inference. Key methods include Quantization-Aware Training (QAT) (Bhalgat et al., 2020) and the more efficient Post-Training Quantization (PTQ) (Li et al., 2021; Nagel et al., 2020), which doesn’t require retraining. For diffusion models, PTQ methods such as PTQ4DiT (Wu et al., 2024) and Q-diffusion (Li et al., 2023b) optimize activation variance across denoising steps, while TDQ (So et al., 2024) uses MLP layers for step-wise parameter estimation. These approaches aim to balance performance with efficiency.
Sampling Optimization. Sampling method optimization for diffusion models aims to reduce generation time while maintaining sample quality. These can be categorized into re-training methods, like knowledge distillation (Luhman and Luhman, 2021; Salimans and Ho, 2022), and those that do not require re-training, such as advanced samplers for pre-trained models. Notably, DDIM (Song et al., 2020) reduces sampling steps using non-Markovian processes, while DPM-Solver (Lu et al., 2022) leverages ODE solvers. Progressive distillation (Meng et al., 2023; Yin et al., 2024), consistency models (Song et al., 2023), and parallel sampling methods like DSNO (Zheng et al., 2023) and ParaDiGMS (Shih et al., 2024) also enhance efficiency without additional training, making diffusion models more practical for real-world applications.
@@ 2.2. Caching Method
Next, Fora (Selvaraju et al., 2024) implemented a caching mechanism that stores and reuses intermediate outputs of attention and MLP layers across denoising steps, reducing computational overhead. $\Delta$-DiT (Chen et al., 2024) uses a designed caching mechanism to accelerate the later DiT blocks in the early sampling stages and the earlier DiT blocks in the later stages. These works extend caching to DiT-based models. PAB (Zhao et al., 2024) reduces redundancy by utilizing a U-shaped attention pattern in the diffusion process and expands caching to video generation through pyramid broadcasting. ToCa (Zou et al., 2024) extends the caching dimension from the block level to the token level, thereby increasing the upper limit of caching acceleration. Beyond the rule-based methods mentioned, learning-to-cache (Ma et al., 2024b) achieves higher acceleration by using a learnable router to decide whether each layer needs computation, though it incurs significant computational costs.
@@ 3.2. Error-Optimized Cache
Next, through observation, the absolute values of each pixel value in the trend vary. For relatively large ${v}_{t}^{l}$, we believe that they are the accumulation of many numbers with the same sign (positive or negative). Therefore, large numbers are representative of the trend and can be used to compensate for the errors caused by caching and reduce the accumulation of cache errors. As shown in Figure 4, the $index$ represents that cache optimization is applied in the first $index$ steps. The dog’s tongue is repaired as the index increases. For relatively small ${v}_{t}^{l}$, we consider it to be an accumulation of positive and negative values around zero. If they are applied in the later stage of generation, they will cause problems such as blurred background details. As shown in Figure 4, when $index=19$, the details of the water in the bird’s background and the grass in the dog’s background are greatly weakened. Therefore, the step at which cache optimization is applied should be in the early stage of generation. In this way, through model operations, the noise with small absolute values will be mapped into the required generated content, thus avoiding the above-mentioned problems.

### v3

@@ Abstract.
Diffusion Transformer (DiT) is a crucial method for content generation. However, it needs a lot of time to sample. Many studies have attempted to use caching to reduce the time consumption of sampling. Existing caching methods accelerate generation by reusing DiT features from the previous time step and skipping calculations in the next, but they tend to locate and cache low-error modules without focusing on reducing caching-induced errors, resulting in a sharp decline in generated content quality when increasing caching intensity. To solve this problem, we propose the Error-Optimized Cache (EOC). This method introduces three key improvements: (1) Prior knowledge extraction: Extract and process the caching differences; (2) A judgment method for cache optimization: Determine whether certain caching steps need to be optimized; (3) Cache optimization: reduce caching errors. Experiments show that this algorithm significantly reduces the error accumulation caused by caching, especially excessive caching. On the ImageNet dataset, without substantially increasing the computational load, this method improves the FID↓ of the generated images when the rule-based model FORA has a caching level of 75%, 50%, and 25%, and the training-based model Learning-to-cache has a caching level of 22%. Specifically, the FID↓ values change from 30.454 to 21.690 (28.8%), from 6.857 to 5.821 (15.1%), from 3.870 to 3.692 (4.6%), and from 3.539 to 3.451 (2.5%) respectively. Code is available at https://github.com/qiujx0520/EOC_MM2025.git.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/XXXXXX.XXXXXXccs: Computing methodologies Computer vision
@@ 1. Introduction
Content generation aims to automatically produce text (Achiam et al., 2023; Lu et al., 2025; Zhu et al., 2024c; Li et al., 2025), images (Xu et al., 2024; Wang et al., 2025; Lu et al., 2023; Zhu et al., 2024a, b), videos (Wang et al., 2024; Qiu et al., 2025; Wang et al., 2018; Guo et al., 2019), and other forms of content based on existing data or context. It is widely used in intelligent customer service dialogue generation, news summarization, and advertising copy creation. However, current generation methods often rely on complex models, resulting in cumbersome and time-consuming computational processes that hinder rapid deployment and flexible application in real-world business contexts.
@@ 2.1. Traditional Generation Acceleration
Model Pruning And Quantization. Pruning reduces model complexity and computational cost by removing less critical weights, neurons, or layers while maintaining performance. It involves unstructured pruning (Dong et al., 2017; Lee et al., 2019), which masks individual parameters, and structural pruning (Liu et al., 2021), which removes larger components like layers or filters. In diffusion models, techniques like LD-Pruner (Castells et al., 2024) and DiffPruning (Guo et al., 2020) enhance efficiency by applying structured pruning and using informative gradients to reduce unnecessary steps. Quantization lowers the precision of weights and activations to smaller bit formats, reducing model size and speeding up inference. Key methods include Quantization-Aware Training (QAT) (Bhalgat et al., 2020) and the more efficient Post-Training Quantization (PTQ) (Li et al., 2021; Nagel et al., 2020), which doesn’t require retraining. For diffusion models, PTQ methods such as PTQ4DiT (Wu et al., 2024) and Q-diffusion (Li et al., 2023b) optimize activation variance across denoising steps, while TDQ (So et al., 2024) uses MLP layers for step-wise parameter estimation. MPQ-DM(Feng et al., 2025) achieves significant accuracy gains at extremely low bit-widths through two techniques: outlier-driven mixed quantization (OMQ) and temporal smooth relationship distillation (TRD). These approaches aim to balance performance with efficiency.
Sampling Optimization. Sampling method optimization for diffusion models aims to reduce generation time while maintaining sample quality. These can be categorized into re-training methods, like knowledge distillation (Luhman and Luhman, 2021; Salimans and Ho, 2022; Feng et al., 2024), and those that do not require re-training, such as advanced samplers for pre-trained models. Notably, DDIM (Song et al., 2020) reduces sampling steps using non-Markovian processes, while DPM-Solver (Lu et al., 2022) leverages ODE solvers. Progressive distillation (Meng et al., 2023; Yin et al., 2024), consistency models (Song et al., 2023), and parallel sampling methods like DSNO (Zheng et al., 2023) and ParaDiGMS (Shih et al., 2024) also enhance efficiency without additional training, making diffusion models more practical for real-world applications.
@@ 2.2. Caching Method
Next, Fora (Selvaraju et al., 2024) implemented a caching mechanism that stores and reuses intermediate outputs of attention and MLP layers across denoising steps, reducing computational overhead. $\Delta$-DiT (Chen et al., 2024) uses a designed caching mechanism to accelerate the later DiT blocks in the early sampling stages and the earlier DiT blocks in the later stages. These works extend caching to DiT-based models. ToCa (Zou et al., 2024) extends the caching dimension from the block level to the token level, thereby increasing the upper limit of caching acceleration. Beyond the rule-based methods mentioned, learning-to-cache (Ma et al., 2024b) achieves higher acceleration by using a learnable router to decide whether each layer needs computation, though it incurs significant computational costs. Building on image generation, caching can also be applied to video generation acceleration(Zhao et al., 2024; Lv et al., 2024; Liu et al., 2025).
@@ 3.2. Error-Optimized Cache
Next, through observation, the absolute values of each pixel value in the trend vary. For relatively large ${v}_{t}^{l}$, we believe that they are the accumulation of many numbers with the same sign (positive or negative). Therefore, large numbers are representative of the trend and can be used to compensate for the errors caused by caching and reduce the accumulation of cache errors. As shown in Figure 4, the $index$ represents that cache optimization is applied in the first $index$ steps. The dog’s tongue is repaired as the index increases. For relatively small ${v}_{t}^{l}$, we consider it to be an accumulation of positive and negative values around zero. If they are applied in the later stage of generation, they will cause problems such as blurred background details. As shown in Figure 4, when $index=19$, the details of the water in the bird’s background and the grass in the dog’s background are greatly weakened. In other words, even if trends are introduced in the early stage of the sampling process, causing errors, the model can map these errors back to the normal range through subsequent sampling steps. However, errors caused by introducing trends in the later stage do not have enough time to return to normal. Therefore, introducing trends in the later stage will result in more negative impacts. Therefore, the step at which cache optimization is applied should be in the early stage of generation. In this way, through model operations, the noise with small absolute values will be mapped into the required generated content, thus avoiding the above-mentioned problems.
@@ 5. Conclusion
Acknowledgements
This research is supported by the Local Science and Technology Program (No.2024CSJGG00800).